# Robust phase estimation of the ground-state energy without controlled time evolution on a quantum device

### v1

@@ Abstract
$\hat{H}_{\mathrm{D}}=\sum_{i=1}^{N/2}J_{2i-1,2i}(\hat{\sigma}^{x}_{2i-1}\hat{\sigma}^{x}_{2i}+\hat{\sigma}^{y}_{2i-1}\hat{\sigma}^{y}_{2i})+\sum_{i=1}^{N}B_{i}\hat{\sigma}^{z}_{i}$ (7)
@@
,where $J_{2i-1,2i}$ is the interaction strength between qubits, $B_{i}$ denotes the strength of the magnetic field acting on the $i$-th qubit. The parameters are set as $J_{1,2}/J=0.5$, $J_{3,4}/J=0.3$, $B_{1}/J=B_{2}/J=-1$, and $B_{3}/J=B_{4}/J=1$. Under these settings, the ground state of the driving Hamiltonian is |$\ket*{1100}$. Here, we consider the case where $\hat{Q}=\hat{M}$, which clearly commutes with both $\hat{H}_{\mathrm{D}}$ and $\hat{H}_{\mathrm{P}}$. Since the states |$\ket*{11\cdots 1}$ and |$\ket*{00\cdots 0}$ are the unique eigenstates of $\hat{M}$ corresponding to the eigenvalues ±$\pm N$, they are trivial simultaneous eigenstates of both $\hat{H}_{\mathrm{D}}$ and $\hat{H}_{\mathrm{P}}$, with their corresponding trivial eigenvalues denoted as $E_{\pm N,\mathrm{D}}$ and $E_{\pm N,\mathrm{P}}$, respectively. The initial state is prepared as follows: |$\ket*{\Psi_{0}}=\frac{1}{\sqrt{2}}(\ket*{1100}+\ket*{1111})$. Our proposed method is implemented starting from this state.
@@
Figure 1 (b) shows the Fourier transform |$\absolutevalue{f(\omega)}$ | plotted against $\omega$. We set the execution time for ASP as $JT=5$ and varied $\tau$ from $J\tau_{\mathrm{min}}=0$ to $J\tau_{\mathrm{max}}=70$ in $L=1000$ steps. The peaks observed at $\omega/J\simeq 10.8,8.1,3.8,2.5$ correspond to the energy differences between the estimated energy eigenvalues and the trivial energy eigenvalue. In the case of the Heisenberg model, for the above parameters, the reference state |$\ket*{1111}$ corresponds to the maximum energy eigenvalue. Due to this, we can identify the ground-state energy from the largest energy difference. Specifically, the peak at $\omega/J=10.8$ corresponds to the energy difference between the trivial eigenvalue $E_{N,\mathrm{P}}$ and the ground-state energy $E^{\mathrm{g}}_{0,\mathrm{P}}$ of the problem Hamiltonian. Given that $E_{N,\mathrm{P}}/J=4.3$, the ground-state energy is estimated as $E^{\mathrm{g}}_{0,\mathrm{P}}/J\simeq-6.524590$. In comparison, the ground-state energy obtained using numerical diagonalization is $E^{\mathrm{g}}_{0,\mathrm{P}}/J\simeq-6.524593$, corresponding to a relative error of $3.7\times 10^{-5}\%$. Furthermore, the energy levels of the excited states obtained via numerical diagonalization, all expressed in units of $J$, are approximately $-3.80585,0.53380$ and $1.849$ for the second, eighth, and eleventh excited states, respectively. The corresponding values estimated from the peaks of $f(\omega)$ are around $-3.80588$, $0.53388$ , and $1.845$, respectively. The relative errors between the exact and estimated values are $6\times 10^{-4}\%$, $1.4\times 10^{-2}\%$, and $2.3\times 10^{-1}\%$, respectively. As the eigenenergy of the excited state increases, the relative error of the estimation also increases. This can be explained as follows. The trivial energy eigenvalue is the largest among the eigenenergies of the problem Hamiltonian. Consequently, the energy gap between the reference state and higher excited states becomes smaller. To estimate the small energy gap, a longer time $\tau$ is required for the Ramsey-type measurement due to the uncertainty relation between time and energy. To achieve high accuracy in estimating the excitation energies, the initial state should be prepared as a superposition of the reference state and the excited state, and then the proposed method can be applied. Alternatively, changing the reference state or choosing a sufficiently long $\tau$ can also resolve this limitation.
@@
Conclusion— In this study, we propose a method to accurately estimate the ground-state energy of a Hamiltonian using ASP on the early FTQC or the quantum annealing, without requiring controlled time evolution. Our approach leverages the symmetry of the Hamiltonian, enabling the direct estimation of the ground-state energy by preparing an initial state that is a superposition of the ground state and a reference state. Unlike the conventional method [29, 30] to estimate the energy gap without controlled time evolution, our method directly estimates the ground-state energy eigenvalue. Comparison with the result obtained by the conventional ASP reveals that our method has the advantage of robustness against non-adiabatic transitions, enabling accurate ground-state energy estimation even at short execution times. Furthermore, it can estimate the energies of excited states induced by non-adiabatic transitions. Future work includes further experimental verification of this method, exploration of its applicability to other quantum systems, and generalization of the symmetry utilized in the initial state preparation.
@@
Acknowledgment— We thank Suguru Endo for his useful comments. This work is supported by JSPS KAKENHI (Grant Number 23H04390), JST Moonshot (Grant Number JPMJMS226C), CREST (JPMJCR23I5), and presto JST. K.S. acknowledges support from Quantum Leap Flagship Program (Grant No. JPMXS0120319794) from the MEXT, Japan, Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from JST, Japan, and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819).

### v2

@@ III Result
$\hat{H}_{\mathrm{D}}=\sum_{i=1}^{N/2}J_{2i-1,2i}(\hat{\sigma}^{x}_{2i-1}\hat{\sigma}^{x}_{2i}+\hat{\sigma}^{y}_{2i-1}\hat{\sigma}^{y}_{2i})+\sum_{i=1}^{N}B_{i}\hat{\sigma}^{z}_{i},$ (14)
@@
where $J_{2i-1,2i}$ is the interaction strength between qubits, $B_{i}$ denotes the strength of the magnetic field acting on the $i$-th qubit. The parameters are set as $J_{1,2}/J=0.5$, $J_{3,4}/J=0.3$, $B_{1}/J=B_{2}/J=-1$, and $B_{3}/J=B_{4}/J=1$. Under these settings, the ground state of the driving Hamiltonian is |$\ket*{1100}$. Here, we consider the case where $\hat{Q}=\hat{M}$, which clearly commutes with both $\hat{H}_{\mathrm{D}}$ and $\hat{H}_{\mathrm{P}}$. Since the states |$\ket*{11\cdots 1}$ and |$\ket*{00\cdots 0}$ are the unique eigenstates of $\hat{M}$ corresponding to the eigenvalues ±$\pm N$, they are trivial simultaneous eigenstates of both $\hat{H}_{\mathrm{D}}$ and $\hat{H}_{\mathrm{P}}$, with their corresponding trivial eigenvalues denoted as $E_{\pm N,\mathrm{D}}$ and $E_{\pm N,\mathrm{P}}$, respectively. The initial state is prepared as follows: |$\ket*{\Psi_{0}}=\frac{1}{\sqrt{2}}(\ket*{1100}+\ket*{1111})$ (see Appendix B for details). Our proposed method is implemented starting from this state.
@@
Figure 1 (b) shows the Fourier transform |$\absolutevalue{f(\omega/J)}$ | plotted against $\omega/J$. We set the execution time for ASP as $JT=5$ and varied $\tau$ from $J\tau_{\mathrm{min}}=0$ to $J\tau_{\mathrm{max}}=70$ in $L=1000$ steps. The peaks observed at $\omega/J\simeq 10.8,8.1,5.4,3.8,2.5$ correspond to the energy differences between the estimated energy eigenvalues and the trivial energy eigenvalue. In the case of the Heisenberg model, for the above parameters, the reference state |$\ket*{1111}$ corresponds to the maximum energy eigenvalue. Due to this, we can identify the ground-state energy from the largest energy difference. Specifically, the peak at $\omega/J=10.8$ corresponds to the energy difference between the trivial eigenvalue $E_{N,\mathrm{P}}$ and the ground-state energy $E^{\mathrm{g}}_{0,\mathrm{P}}$ of the problem Hamiltonian. Given that $E_{N,\mathrm{P}}/J=4.3$, the ground-state energy is estimated as $E^{\mathrm{g}}_{0,\mathrm{P}}/J\simeq-6.524590$. In comparison, the ground-state energy obtained using numerical diagonalization is $E^{\mathrm{g}}_{0,\mathrm{P}}/J\simeq-6.524593$, corresponding to a relative error of $3.7\times 10^{-5}\%$. Furthermore, the energy levels of the excited states obtained via numerical diagonalization, all expressed in units of $J$, are approximately $-3.80585,-1.09182,0.53380$ and $1.84517$ for the second, fifth, seventh, and tenth excited states, respectively. The corresponding values estimated from the peaks of $f(\omega)$ are around $-3.80588$, $-1.08495$, $0.53388$ , and $1.84942$, respectively. The relative errors between the exact and estimated values are $6.0\times 10^{-4}\%$, $6.0\times 10^{-1}\%$, $1.4\times 10^{-2}\%$, and $2.3\times 10^{-1}\%$, respectively.
There is a tendency for the relative error of the estimation to increase as the eigenenergy of the excited state increases. A possible explanation is as follows. The trivial energy eigenvalue is the largest among the eigenenergies of the problem Hamiltonian. Consequently, the energy gap between the reference state and higher excited states becomes smaller. To estimate this small energy gap, a longer time $\tau$ is required for the Ramsey-type measurement due to the time-energy uncertainty relation. To achieve high accuracy in estimating the excitation energies, the initial state should be prepared as a superposition of the reference state and the excited state of the driving Hamiltonian, and then the proposed method can be applied. Alternatively, changing the reference state or choosing a sufficiently long $\tau$ can also overcome the challenge of accurately estimating the eigenenergy of the excited state.
@@ IV Conclusion
In this study, we propose a method to accurately estimate the ground-state energy of a Hamiltonian using ASP in early FTQC or quantum annealing, without requiring controlled time evolution. Our approach leverages the symmetry of the Hamiltonian, enabling the direct estimation of the ground-state energy by preparing an initial state that is a superposition of the ground state and a reference state. Unlike the conventional method [29, 30] for estimating the energy gap without controlled time evolution, our method directly estimates the ground-state energy eigenvalue in the subspace of interest. To identify the true ground-state energy of the problem Hamiltonian from the full spectrum, a search across all subspaces can be performed in polynomial time, owing to the second assumption of our protocol (see Appendix C for details). Furthermore, a comparison with the result obtained by the conventional ASP reveals that our method offers a clear practical advantage. It is robust against non-adiabatic transitions, eliminating the need for fine-tuning the execution time $T$ and enabling accurate ground-state energy estimation even with short execution times. This is in stark contrast to the conventional ASP, which relies on strict adiabaticity. Moreover, it can estimate the energies of excited states induced by non-adiabatic transitions. Future work includes further experimental verification of this method, exploration of its applicability to other quantum systems, and generalization of the symmetry utilized in the initial state preparation.
V Acknowledgment
@@
We thank Suguru Endo for his useful comments. This work is supported by JSPS KAKENHI (Grant Number 23H04390, 20H05661), JST Moonshot (Grant Number JPMJMS226C), CREST (JPMJCR23I5), and presto JST (JPMJPR245B). K.S. acknowledges support from Quantum Leap Flagship Program (Grant No. JPMXS0120319794) from the MEXT, Japan, Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from JST, Japan, and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819).
Appendix A Effect of relative phases on phase estimation
In this appendix, we provide a mathematical proof that the relative phase factor $e^{i\theta}$ of the superposition state |$\ket{\psi}$ between different subspaces used in our method does not affect the final eigenvalue estimation, regardless of its specific value.
Let us rewrite $\mathcal{P}(\tau)$ given by Eq. (10) as
$\mathcal{P}(\tau)=\cos^{2}\quantity(\frac{\Delta E\tau}{2}+\frac{\theta^{\prime}}{2})=\frac{1}{2}+\frac{1}{2}\cos\quantity(\frac{\Delta E\tau}{2}+\frac{\theta^{\prime}}{2}),$ (15)
where we define $\Delta E\equiv E^{\mathrm{g}}_{q^{\prime},\mathrm{P}}-E^{n}_{q,\mathrm{P}}$. Substituting this into Eq. (11), we obtain
$f(\omega)=\frac{1}{2}\sum_{n=1}^{L}e^{-i\omega\tau_{n}}+\frac{1}{2}\sum_{n=1}^{L}\cos\quantity(\frac{\Delta E\tau_{n}}{2}+\frac{\theta^{\prime}}{2})e^{-i\omega\tau_{n}}.$ (16)
[…]
The key observation here is that the relative phase $\theta^{\prime}$ factors out as a global phase term. The Fourier transform determines the peak positions and magnitudes based on the term $\exp\quantity[i\frac{(\Delta E-2\omega)\tau_{n}}{2}]$, while the phase factor $\exp\quantity(i\frac{\theta^{\prime}}{2})$ has an absolute value of $1$ and thus does not affect the final eigenvalue estimation.
Appendix B Initial state preparation via ASP
An implicit assumption of our proposal, as described in the main text, is that the initial state can be prepared as a superposition of the ground state (belonging to a subspace characterized by a specific eigenvalue $q$) of the driving Hamiltonian and the reference state. In this appendix, we discuss how such a state in the case of $\hat{Q}=\hat{M}$ can be prepared. The driving Hamiltonian used in this study is given by Eq. (14). The initial state is defined as
|$\ket{\psi_{\mathrm{ini}}}=\frac{1}{\sqrt{2}}\quantity(\ket{\psi^{\mathrm{g}}_{m,\mathrm{D}}}+\ket{11\cdots 1}),$ (18)
where |$\ket{\psi^{\mathrm{g}}_{m,\mathrm{D}}}$ represents the ground state of the driving Hamiltonian within the subspace corresponding to each eigenvalue $m$ of the conserved quantity $\hat{M}$, and |$\ket{11\cdots 1}$ represents the reference state.
To begin with, the driving Hamiltonian in Eq. (14) should be diagonalized to identify the energy levels of |$\ket{\psi^{\mathrm{g}}_{m,\mathrm{D}}}$ and the reference state |$\ket{11\cdots 1}$ within the full energy spectrum. Specifically, if diagonalization reveals that these states correspond to the $n$-th and $l$-th excited states,respectively, we can use this information to generate the desired initial state by using ASP.
The basic strategy is as follows. The desired state is $\frac{1}{\sqrt{2}}(|\tilde{E}_{n,\mathrm{D}}\rangle+|\tilde{E}_{l,\mathrm{D}}\rangle)$ where $|\tilde{E}_{n,\mathrm{D}}\rangle$ ($|\tilde{E}_{l,\mathrm{D}}\rangle$) is the $n$-th ($l$-th) eigenstate of the driving Hamiltonian in Eq. (14) Once we create $\frac{1}{\sqrt{2}}(|\tilde{E}_{n}\rangle+|\tilde{E}_{l}\rangle)$ where $|\tilde{E}_{n}\rangle$ ($|\tilde{E}_{l}\rangle$) is the $n$-th ($l$-th) eigenstate of the Hamiltonian of the transverse magnetic field $\hat{H}_{\mathrm{transverse}}$, we can adiabatically prepare the desired state by changing the Hamiltonian from $\hat{H}_{\mathrm{transverse}}$ to $\hat{H}_{\mathrm{D}}$. Although it is known that we can prepare the so-called GHZ state, $\frac{1}{\sqrt{2}}(|\tilde{E}_{n}\rangle+|\tilde{E}_{l}\rangle)$ is slightly different from the GHZ state. Fortunately, after creating a small GHZ state, we can prepare just by adding separable qubits, $\frac{1}{\sqrt{2}}(|\tilde{E}_{n}\rangle+|\tilde{E}_{l}\rangle)$. By combining these strategy, our proposed prescription is as follows.
For simplicity, let us assume that the number of qubits are even. The procedure consists of the following steps:
  1. 1.
    The process starts from the all-plus state $\ket{++\cdots+}$. We use ASP to transform this state into a GHZ state $|\psi_{\rm{GHZ}}\rangle=(\ket{00\cdots 0}+\ket{11\cdots 1})\sqrt{2}$ by gradually switching the Hamiltonian from $\hat{H}_{1}=B_{1}\sum_{i}\hat{\sigma}^{x}_{i}$ to $\hat{H}_{2}=-B_{2}{(\sum_{i}\hat{\sigma^{z}_{i}})}^{2}$, where $B_{1}$ ($B_{2}$) represent the strength of the magnetic field (interaction between qubits) [45, 46, 47, 48, 49, 50, 51].
  2. 2.
    The Hamiltonian is then switched from $\hat{H}_{2}$ to $\hat{H}_{3}=-B_{3}{(\sum_{i}\hat{\sigma^{x}_{i}})}^{2}$ by using ASP, where $B_{3}$ represent the strength of the interaction between qubits, so that the state becomes $|\psi^{\prime}_{\rm{GHZ}}\rangle=(\ket{++\cdots+}+\ket{--\cdots-})\sqrt{2}$. Before and after this ASP step, the system remains a superposition of degenerate ground states. Both states $|\psi_{\rm{GHZ}}\rangle$ and $|\psi^{\prime}_{\rm{GHZ}}\rangle$ are eigenstates of a parity operator defined as $\hat{P}=\hat{\sigma}^{x}_{1}\otimes\hat{\sigma}^{x}_{2}\otimes\cdots\otimes\hat{\sigma}^{x}_{N}$.
  3. 3.
    The Hamiltonian is switched from $\hat{H}_{3}$ to $\hat{H}_{4}=\sum_{i}B_{i}\hat{\sigma}^{x}_{i}$. Note that, because $\hat{H}_{3}$ and $\hat{H}_{4}$ commute, the state remains unchanged.
  4. 4.
    By adding qubits to match the desired superposition, we can create a state such as |±±±(|++++|$\ket{\pm\pm\cdots\pm}(\ket{++\cdots+}+\ket{--\cdots-})/\sqrt{2}$. After relabeling the qubits, this corresponds to $\frac{1}{\sqrt{2}}(|\tilde{E}_{n}\rangle+|\tilde{E}_{l}\rangle)$ where $|\tilde{E}_{n}\rangle$ ($|\tilde{E}_{l}\rangle$) is the $n$-th ($l$-th) eigenstate of $\hat{H}_{\mathrm{transverse}}=\sum_{i}B_{i}\hat{\sigma}^{x}_{i}$, where $\ket{\pm\pm\cdots\pm}$ represents a product state in which each qubit is in either $\ket{+}$ or $\ket{-}$.
  5. 5.
    Finally, the Hamiltonian is switched from $\hat{H}_{\mathrm{transverse}}$ to $\hat{H}_{\mathrm{D}}$ (Eq. (14)) by using ASP. Then, we obtain the desired superposition state (Eq. (18)).
At the final step, $\hat{H}_{\mathrm{transverse}}$ is the Hamiltonian without interaction while $\hat{H}_{\mathrm{D}}$ is the Hamiltonian with interaction only between the $12i-$ 1-th qubit and the $2i$-th qubit. This means that the evolution at the final step can be reduced to two-qubit dynamics. It is worth mentioning that we can reduce the number of qubits by measuring one of the qubits of the GHZ states with an even number of qubits if we need a GHZ state with an odd number of qubits.
Appendix C Phase estimation in different subspaces
Figure 2: The amplitude $|f(\omega/J)|$ obtained from the numerical simulation as a function of the frequency $\omega/J$. (a) Results for a $m=2$ subspace. (b) A magnified view of a specific region in (a). (c) Results for a $m=-2$ subspace. (d) A magnified view of a specific region in (c). (a),(b) The peaks were observed at $\omega/J\simeq 8.6,6.0,3.3$, and $1.9$. These peaks correspond to the energy differences between the reference state of $\hat{H}_{\mathrm{P}}$ and the first, fourth, eighth, and eleventh excited states, respectively. (c),(d) The peaks were observed at $\omega/J\simeq 7.7$ and $4.7$. These peaks correspond to the energy differences between the reference state of $\hat{H}_{\mathrm{P}}$ and the third, and sixth excited states, respectively. Note that the peaks not discussed here correspond to the energy differences between the excited states.
In this study, we propose a method for estimating the energy eigenvalues within a subspace characterized by a specific eigenvalue $q$ of a conserved quantity $\hat{Q}$, which commutes with both the driving Hamiltonian and the problem Hamiltonian. An key aspect of our method is that it does not require prior knowledge of which subspace contains the true ground state of the problem Hamiltonian. By applying our method to each subspace separately and comparing the obtained eigenvalues, we can identify the true ground state of the problem Hamiltonian. Of course, if the subspace containing the target eigenvalue of the problem Hamiltonian is known in advance through some other method, our approach can be directly applied to that subspace to obtain the desired eigenvalue. Alternatively, if the ground state of a particular subspace is of interest, our method can also be used to identify it.
In Sec. III, we considered the case of $\hat{Q}=\hat{M}$ and applied our proposed method to the subspace with eigenvalue $m=0$ to estimate the ground energy. In the main text, we made this choice for simplicity assuming that prior knowledge from exact diagonalization indicated that the true ground state of Eq. (13) lies in the $m=0$ subspace. In this appendix, we remove the assumption. More specifically, we apply our method to the subspaces with $m=\pm 2$ and verify that the true ground energy can be estimated from the obtained eigenvalues. However, we omit the cases of $m=\pm 4$, as these subspaces are both one-dimensional and therefore trivial.
Figure 2 (a) and (b) show the Fourier transform |$\absolutevalue{f(\omega/J)}$ | plotted against $\omega/J$ for $m=2$. The numerical parameters used are the same as those in Sec. III, except for the initial states. For the $m=2$ subspace, we set the initial state as a superposition of |$\ket*{\psi^{\mathrm{g}}_{2,\mathrm{D}}}=\frac{1}{\sqrt{2}}(\ket{0010}-\ket{0001})$ and the reference state, which was chosen as |$\ket{1111}$. The results show peaks at $\omega/J\simeq 8.6,6.0,3.3$, and $1.9$. These peaks correspond to the energy differences between the estimated energy eigenvalues and the trivial energy eigenvalue. Given that $E_{4,\mathrm{P}}/J=4.3$, the estimated energy eigenvalues from the peaks of |$\absolutevalue{f(\omega/J)}$ | are approximately $-4.28497,-1.68560,0.97414$, and $2.38026$, all expressed in units of $J$. The energy levels of the excited states obtained via numerical diagonalization, all expressed in units of $J$, are approximately $-4.28498,-1.68559,0.97265$ and $2.37795$ for the first, fourth, eighth, and eleventh excited states, respectively. Similarly, we plot the results for $m=-2$ in Fig. 2 (c) and (d). Here, the initial state is a superposition of |$\ket*{\psi^{\mathrm{g}}_{-2,\mathrm{D}}}=\frac{1}{\sqrt{2}}(\ket{1011}-\ket{0111})$ and the reference state |$\ket{1111}$.
The results in Fig. 2(c) and (d) show peaks at $\omega/J\simeq 7.7$ and $4.7$. The estimated energy eigenvalues from the peaks of |$\absolutevalue{f(\omega/J)}$ | are approximately $-3.41947$ and $-0.36850$. The energy levels of the excited states obtained via numerical diagonalization, all expressed in units of $J$, are approximately $-3.41949$ and $-0.36855$ for the third and sixth excited states, respectively. Note that the peaks not discussed here correspond to the energy differences between the excited states. By comparing these results for $m=\pm 2$ with those for $m=0$, we can conclude that the energy $-6.524590$ mentioned in Sec. III is indeed the true ground-state energy of the problem Hamiltonian.
As these results indicate, our proposed method does not require prior knowledge of which subspace contains the true ground energy of the problem Hamiltonian. By applying our method to different subspaces and comparing the obtained results, we can successfully estimate the true ground energy of the problem Hamiltonian.